# Combining Flow Matching and Transformers for Efficient Solution of Bayesian Inverse Problems

### v1

@@ Abstract
Solving Bayesian inverse problems efficiently remains a significant challenge due to the complexity of posterior distributions and the computational cost of traditional sampling methods. Given a series of observations and the forward model, we want to recover the distribution of the parameters, conditioned on observed experimental data. We show, that combining Conditional Flow Mathching (CFM) with transformer-based architecture, we can efficiently sample from such kind of distribution, conditioned on variable number of observations.
Keywords Conditional Flow Matching  $\cdot$ Bayesian Inverse Problems  $\cdot$ Generative Modeling  $\cdot$ Inverse Problems
@@ 1 Introduction
Many natural processes can be mathematically modeled using appropriate formal representations. However, the challenge often lies in inferring latent parameters that are not directly observable. These parameters must typically be estimated from limited observations, giving rise to Bayesian inverse problems. The idea of Bayesian inversion is to parametrize the posterior distribution of model parameters, given observations and a prior distribution on the model parameters. The main challenge is that typically the distribution is known up to a normalization constant, making sampling from the posterior intractable. Classical methods like Markov Chain Monte Carlo (MCMC) Geyer [1992] rely on many forward problem solutions for each set of observations, which can be very time-consuming.
@@
  • We formulate the Bayseian inverse problem as the problem of learning conditional probability distribution from samples, that can be easily constructed
  • We propose a transformer-based CFM architecture that can handle different number of observations
@@ 2 Background and Related Work
Classical approaches to solve Bayesian Inverse Problem
@@
Bayesian inverse problems have also been addressed using physics-informed neural networks [Raissi et al., 2019] with combining invertible flow-based neural networks (INN). In [Guan et al., 2023], this approach is shown to be effective but not universally applicable, as it requires designing a custom loss function for each PDE to ensure efficient training. Additionally, the number of observations for the Darcy flow problem remains fixed.
A significant number of inverse problems in medicine can be effectively addressed using generative networks [Aali et al., 2023, Song et al., 2022].
@@
GANs have also been integrated into hybrid Bayesian frameworks. In Patel et al. [2020], a GAN was employed to approximate high-dimensional parameter priors within MCMC sampling. Similarly, Xia and Zabaras [2022] combined a VAE-based prior with MCMC for posterior estimation. This method was called Multiscale deep generative model (MDGM). Although, these methods leverage generative models to enhance prior representation, their computational gains remain limited, as they still require iterative forward model evaluations during sampling. Other works, such as Goh et al. [2021] using VAEs embed generative models into variational inference frameworks. However, these approaches face trade-offs between approximation accuracy and scalability in high-dimensional settings.
However these approaches have two fundamental disadvantages. First, they do not operate with an arbitrary number of observations, which can be critical in real-world problems. Second, these approaches often do not escape the iterative process itself due to the fact that such approaches do not explicitly generate a posterior distribution.
@@
Flow Matching (FM) [Lipman et al., 2023] is an efficitient approach to generative modeling based on Continuous Normalizing Flows (CNFs) [Chen et al., 2018], enabling large-scale CNF training without simulation by regressing vector fields of fixed conditional probability paths. It generalizes diffusion models by supporting a broader class of Gaussian probability paths, including Optimal Transport (OT) displacement interpolation, which enhances efficiency, stability, and sample quality. Compared to diffusion-based methods, FM allows faster training and sampling, improves likelihood estimation, and enables reliable sample generation using standard numerical ODE solvers, making it a alternative for high-performance generative modeling.
@@
Table 1: Comparison of methods for solving Bayeisan Inverse problems. *MDGM use the PDE solution as a holistic observation; the problem was not formulated as the recovery of the forward model from a small number of observations
The table 1 provides methods comparing for solving Bayesian inverse problems. Deep learning methods for solving Bayesian inverse problems exhibit distinct strengths and limitations. MDGM [Xia and Zabaras, 2022] leverages a VAE-based convolution neural network with MCMC for multiscale inference, excelling in high-dimensional PDE-based problems but lacking exact likelihood estimation and flexibility for arbitrary observations. MCGAN [Mücke et al., 2022] combines MCMC with GANs for high-fidelity posterior sampling but suffers from computational complexity, lack of explainability, and fixed observation models. PI-INN [Guan et al., 2023] employs physics-informed flow-based models, enabling exact likelihood estimation and end-to-end training but struggles with variable observation sizes due to architectural constraints. In contrast, CFM-Tr integrates conditional flow matching with transformers, offering exact likelihood estimation, end-to-end training, and adaptability to arbitrary observations, making it suitable for dynamic inverse problems like real-time medical imaging. While MDGM and PI-INN are effective for structured problems, CFM-Tr addresses key limitations by combining flexibility, exact inference, and scalability.
@@ 3 Methodology
Input: Dataset of paired samples $(x_{1},e,d)$, neural network model $\mathbf{v}_{\theta}(t,x,e,d)$, conditioning data $e$ and $d$, time $t\sim\text{Uniform}(0,1)$, number of epochs $N_{\text{epoch}}$
Output: Trained conditional flow model $\mathbf{v}_{\theta}(t,x,e,d)$
for $1$ to $N_{\text{epoch}}$ do
      for each minibatch of samples $(x_{0},x_{1})$ do
             $t\sim\mathcal{U}(0,1)$
              // Sample t
            $x_{0}\sim\text{prior distribution}$
            $x_{t}\leftarrow t\cdot x_{1}+(1-t)\cdot x_{0}$
            Compute the target velocity: $u_{t}\leftarrow x_{1}-x_{0}$
            Predict the velocity: $v_{t}\leftarrow\mathbf{v}(t,x_{t},e,d)$
            Compute the loss: $\mathcal{L}(\theta)\leftarrow\mathbb{E}\left[(v_{t}-u_{t})^{2}\right]$
            Compute gradients: $\nabla_{\theta}\mathcal{L}(\theta)$
       end for
end for
return $\mathbf{v}_{\theta}(t,x,e,d)$
Algorithm 1 Conditional Flow Matching Training Algorithm
@@
The key idea is that we can easily sample from the joint distribution $(m_{i},d_{i},e_{i})$. In order to do that. we generate random model parameters (from the prior distribution) and random observation points. When a pair $m_{i},e_{i}$ is given, we can compute $d_{i}$ using the forward model. But $m_{i}$ is also a sample from the conditional distribution $\pi(m|d_{i},e_{i})$. Thus, when a forward model and prior distributions of model parameters $m$ and experimental parameters $e$ are known, we generate training data by sampling multiple variants of $m$ and $e$ and computing the forward model to obtain observations $d$. For each model parameter $m_{i}$ we sample $d_{i}$ for several points $e_{i}$, so our training data contains a set of the form $(m_{i},d_{i},e_{i})$, where $d_{i}$ and $e_{i}$ are sequences of different lengths. The model should be able to sample $m_{i}$ given observations $(d_{i},e_{i})$. In order to do that, we utilize CFM.
@@
We define a conditional interpolation path between $(m_{0},d,e)$ and $(m,d,e)$, where $(d,m,e)$ is sampled from the dataset. The interpolation path is given by:
@@
Figure 1: Solving the inverse problem using flow-matching scheme
@@
We parametrize the velocity field by the transformer architecture with bi-directional attention, motivated by the Diffusion Transformer Peebles and Xie [2022]. Specifically, our transformer implementation uses linear projection of input parameters into the embedding space. Time is encoded using a Timestep Embedder (from [Peebles and Xie, 2022]), which ensures proper time representation in the embedding space. Root Mean Square (RMS) normalization stabilizes learning dynamics. The activation function is $x=\text{ReLU}(x)^{2}$. Self-attention uses rotary position embeddings (RoPE), enabling the transformer to learn relative token positions and generalize to sequences longer than those seen during training.
The architecture varies slightly across tasks to accommodate different input data representations. Specific implementations for tasks from Section 4 are detailed in Figure 6 in Appendix A.
@@
Output: Generated parameters $m$
$x(t=0)\sim\text{prior distribution}$
$x(t=1)\leftarrow\frac{dx}{dt}=\mathbf{v}_{\theta}(t,x_{t},e,d)$
return $x(t=1)$
Algorithm 2 Conditional Flow Matching Inference Algorithm
Figure 2: Transformer architecture
@@
We need to be able to predict the model parameters from different number of observation points. As mentioned before, we generate datasets with varying numbers of observation points, where each batch corresponds to samples with a specific number of points. During training, the model processes batches with different numbers of points sequentially. We propose two strategies: the first follows Algorithm 1, computing gradients for each batch and updating after a fixed number of batches. Alternatively, as shown in Algorithm 3, we can accumulate gradients across batches with different numbers of points before performing a single optimizer step.
@@ 4.1 Simple nonlinear model
$d(e,m)=e^{2}m^{3}+m\exp(-|0.2-e|)+\eta$
  • Sampling $m$ from $\mathcal{U}[0,1]$
  • Sampling $e$ from $\mathcal{U}[0,1]$
  • Sampling noise $\eta$ from the noise distribution
  • Computing $d=f(m,e)+\eta$
@@ 4.2 SEIR disease model
The SEIR (Susceptible-Exposed-Infected-Recovered) model is a mathematical framework used to simulate the spread of infectious diseases. In this case study, we simulate a realistic scenario where we measure the number of infected and deceased individuals at random times and use this information to recover the control parameters of the ODE system.
$\frac{dS}{dt}$ $=-\beta(t)SI,\frac{dE}{dt}=\beta(t)SI-\alpha E$
$\frac{dI}{dt}$ $=\alpha E-\gamma(t)I,\frac{dR}{dt}=\gamma(t)I$
where $S(t)$, $E(t)$, $I(t)$, $R(t)$ denote the fractions of susceptible, exposed, infected and removed individuals at time $t$, respectively. These are initialized with $S(0)=99$, $E(0)=1$, $I(0)=R(0)=0$.
$\beta(t)$ $=\beta_{1}+\frac{\tanh(7(t-\tau))}{2}(\beta_{2}-\beta_{1})$
$\gamma(t)$ $=\gamma^{r}+\gamma^{d}(t)$
We fix $\tau=2.1$ over a time interval of $[0,4]$. The experiment consists of choosing four time points $e=[a_{1},a_{2},a_{3},a_{4}]\sim\mathcal{U}[1,3]$ to measure the number of infected and deceased individuals $d_{i}=[I_{e_{i}},R_{e_{i}}]$ for $i\in[1,4]$ ($d\in\mathbf{R}^{2\times 4}$). The goal is to optimally infer the 6 rates $\mathbf{m}=[\beta_{1},\alpha,\gamma^{r},\gamma^{d}_{1},\beta_{2},\gamma^{d}_{2}]$. After training an MLP and solving the flow matching problem, we learn to smoothly transition from the distribution $\mathcal{U}[0,1]^{6}$ to the distribution $\hat{\mathbf{m}}\sim\rho(\mathbf{m}|\mathbf{e},\mathbf{d})$.
To summarize the inputs and outputs:
$\varepsilon=\frac{\|d-\hat{d}\|_{2}}{\|d\|_{2}}$
@@ 4.3 Permeability field inversion
$-\nabla\cdot(\kappa\nabla u)=0$
with boundary conditions:
$m(x,\mathbf{m})\approx\sum_{i=1}^{n_{m}}m_{i}\sqrt{\lambda_{i}}\phi_{i}(x),$
The transformer architecture enables the use of different input types for the same problem. Here, in addition to the equation values themselves, we use the coordinates of measurement points. The specific architecture is detailed in Figure 6 in Appendix A.
Our results show that we can effectively recover the PDE coefficient using just a few strategically placed measurement points. Figure 5 demonstrates that with 8 relatively uniformly distributed points over the solution field, we can obtain an almost identical solution (approximately 2.75
@@ 5 Results and Discussions
The true solution of the ODE system and the reconstructed parameter distribution, obtained using only 4 data points, are illustrated in Figure 4.
@@
Figure 5: PDE coefficient and solution: true (left) and reconstructed using Flow-matching (right)
@@
We compared our method with the Metropolis-Hastings MCMC (MH-MCMC) algorithm. For a fair comparison, we ran MH-MCMC with enough iterations to achieve approximately the same relative error as our method (shown in Table 2). The results for the SEIR problem are presented in Table 3.
@@
Comparing Tables 2 and 3, we observe that MCMC requires significantly more iterations to achieve similar accuracy as Conditional Flow Matching. This computational cost is particularly significant for elliptic PDEs. Even with 10,000 MCMC iterations, the error in solving the PDE remains above 30% for 6 or more observations, while Conditional Flow Matching achieves 2-8% error in these cases. The computational time difference is substantial: MCMC takes approximately 37 minutes for 10,000 samples, while Flow Matching requires only 1.08 seconds (on CPU) per inference using Algorithm 2 for the permeability inversion problem. For the SEIR model, the CFM transformer takes approximately 0.22 seconds.
Our numerical experiments demonstrate the following key findings:
@@
Additionally, Conditional Flow Matching shows promise for determining optimal experiment design parameters $e$, which could further enhance its applicability to real-world problems.
@@ 6 Limitations and Future Work
The CFM approach is not the only generative modelling technique available. Certainly, we can use normalizing flows, tensor methods Koval et al. [2024], generative adversarial networks. We also need to establish the efficiency of the approach for high-dimensional Bayesian inverse problems, which will require modifications of the architecture and additional scaling of the datasets. CFM learns to sample from the distribution, whereas the Bayesian optimal experiment design requires the evaluation of the log-likelihood. Although it is in principle possible, we did not study the complexity and accuracy of the evaluation of the logarithm of the posterior distribution.
Another line of work is to introduce the guidance to the CFM objective that is enforced by the forward model. Once we got the estimate of the parameter $m$, we can check, if it really fits the observations; a good question is how to modify the inference procedure to correct for possible errors.
An important limitation is the study of the actual properties of the learned distribution. For the cases, when the model parameters are defined by the observations, the velocity field $v_{t}$ will likely not depend on the noise, but only on $d,e$, effectively solving the regression problem of predicting $m_{i}$ from $d_{i},e_{i}$. The usefulness and emergency of randomization with respect to the noise in CFM still needs to be studied. There are two options for using it. First, we can sample different noises, get estimates of the model parameters and plot the distributions (see the Section on the SEIR model, where some of the parameters show greater variability). The second option is to pick samples that provide better fit to the data. This still needs to be studied, since in some of the experiments we found differences in the distribution of provided by MCMC and CFM. The question has to be studied in more details.

### v2

# Bayesian Inverse Problems Meet Flow Matching: Efficient and Flexible Inference via Transformers

@@ Abstract
The efficient resolution of Bayesian inverse problems remains challenging due to the high computational cost of traditional sampling methods. In this paper, we propose a novel framework that integrates Conditional Flow Matching (CFM) with a transformer-based architecture to enable fast and flexible sampling from complex posterior distributions. The proposed methodology involves the direct learning of conditional probability trajectories from the data, leveraging CFM’s ability to bypass iterative simulation and transformers’ capacity to process arbitrary numbers of observations. The efficacy of the proposed framework is demonstrated through its application to three problems: a simple nonlinear model, a disease dynamics framework, and a two-dimensional Darcy flow Partial Differential Equation. The primary outcomes demonstrate that the relative errors in parameters recovery are as low as 1.5%, and that the inference time is reduced by up to 2000 times on CPU in comparison with the Monte Carlo Markov Chain. This framework facilitates the expeditious resolution of Bayesian problems through the utilisation of sampling from the learned conditional distribution.
@@ 1 Introduction
Many natural processes can be mathematically modeled using appropriate formal representations. However, the challenge often lies in inferring latent parameters that are not directly observable. These parameters must typically be estimated from limited observations, giving rise to Bayesian inverse problems. The idea of Bayesian inversion is to parametrize the posterior distribution of model parameters, given observations and a prior distribution on the model parameters. The main challenge is that typically the distribution is known up to a normalization constant, making sampling from the posterior intractable. Classical methods like Markov Chain Monte Carlo (MCMC) [Geyer, 1992] rely on many forward problem solutions for each set of observations, which can be very time-consuming.
@@
  • We formulate the Bayesian inverse problem as the problem of learning the conditional probability distribution from samples, that can be easily constructed.
  • We propose a transformer-based Conditional Flow Matching (CFM) Lipman et al. [2023] architecture that can handle different number of observations.
@@ 2 Background and Related Work
Classical Approaches to Solving Bayesian Inverse Problems
@@
Bayesian inverse problems have also been addressed using physics-informed neural networks [Raissi et al., 2019] by combining invertible flow-based neural networks (INNs). In [Guan et al., 2023], this approach is shown to be effective but not universally applicable, as it requires designing a custom loss function for each PDE to ensure efficient training. Additionally, the number of observations in the Darcy flow problem is fixed.
A significant number of inverse problems in medicine can be effectively addressed using generative networks [Aali et al., 2023, Song et al., 2022]. In Cunningham et al. [2024] proposed Simformer, a framework for simulation-based inference that combines transformer architectures with score-based diffusion models. Simformer relies on stochastic diffusion processes and does not directly learn deterministic transport maps, which can limit interpretability and inference speed in structured inverse problems.
@@
GANs have also been integrated into hybrid Bayesian frameworks. In Patel et al. [2020], a GAN was employed to approximate high-dimensional parameter priors within MCMC sampling. Similarly, Xia and Zabaras [2022] combined a VAE-based prior with MCMC for posterior estimation. This method was called Multiscale deep generative model (MDGM). Although these methods leverage generative models to enhance prior representation, their computational gains remain limited, as they still require iterative forward model evaluations during sampling. Other works, such as Goh et al. [2021] using VAEs embed generative models into variational inference frameworks. However, these approaches face trade-offs between approximation accuracy and scalability in high-dimensional settings.
However these approaches have two fundamental disadvantages. First, they cannot handle an arbitrary number of observations, which can be critical in real-world problems. Second, these approaches often do not escape the iterative process itself due to the fact that such approaches do not explicitly generate a posterior distribution.
@@
Flow Matching (FM) [Lipman et al., 2023] is an efficitient approach to generative modeling based on Continuous Normalizing Flows (CNFs) [Chen et al., 2018], enabling large-scale CNF training without simulation by regressing vector fields of fixed conditional probability paths. It generalizes diffusion models by supporting a broader class of Gaussian probability paths, including Optimal Transport (OT) displacement interpolation, which enhances efficiency, stability, and sample quality. Compared to diffusion-based methods, FM allows faster training and sampling, improves likelihood estimation, and enables reliable sample generation using standard numerical ODE solvers, making it an alternative for high-performance generative modeling.
@@
Table 1: Comparison of methods for solving Bayesian Inverse problems. *MDGM use the PDE solution as a holistic observation; the problem was not formulated as the recovery of the forward model from a small number of observations
Table 1 compares methods for solving Bayesian inverse problems. Deep learning methods for solving Bayesian inverse problems exhibit distinct strengths and limitations. MDGM [Xia and Zabaras, 2022] leverages a VAE-based convolution neural network with MCMC for multiscale inference, excelling in high-dimensional PDE-based problems but lacking exact likelihood estimation and flexibility for arbitrary observations. MCGAN [Mücke et al., 2022] combines MCMC with GANs for high-fidelity posterior sampling but suffers from computational complexity, lack of explainability, and fixed observation models. PI-INN [Guan et al., 2023] employs physics-informed flow-based models, enabling exact likelihood estimation and end-to-end training but struggles with variable observation sizes due to architectural constraints. In contrast, CFM-Tr integrates conditional flow matching with transformers, offering exact likelihood estimation, end-to-end training, and adaptability to arbitrary observations, making it suitable for dynamic inverse problems like real-time medical imaging. While MDGM and PI-INN are effective for structured problems, CFM-Tr addresses key limitations by combining flexibility, exact inference, and scalability.
@@ 3 Methodology
Figure 1: Solving the inverse problem using flow-matching scheme
@@
The key idea is that we can easily sample from the joint distribution $(m_{i},d_{i},e_{i})$. In order to do that. we generate random model parameters (from the prior distribution) and random observation points. When a pair $m_{i},e_{i}$ is given, we can compute $d_{i}$ using the forward model. Importantly, $m_{i}$ is also a sample from the conditional distribution $\pi(m|d_{i},e_{i})$. Thus, when a forward model and prior distributions of model parameters $m$ and experimental parameters $e$ are known, we generate training data by sampling multiple variants of $m$ and $e$ and computing the forward model to obtain observations $d$. For each model parameter $m_{i}$ we sample $d_{i}$ for several points $e_{i}$, thus, our training data consists of tuples of the form of the form $(m_{i},d_{i},e_{i})$, where $d_{i}$ and $e_{i}$ may have variable lengths. The model should be able to sample $m_{i}$ given observations $(d_{i},e_{i})$. In order to do that, we utilize CFM.
@@
We define a conditional interpolation path between $(m_{0},d,e)$ and $(m,d,e)$, where $(d,m,e)$ is drawn from the dataset.
$\mathbb{E}_{t\sim\mathcal{U}(0,1)}\mathbb{E}_{m_{0}\sim\text{prior}}\mathbb{E}_{(m,d,e)\sim data}\left[\big{\|}v_{\theta}(m_{t},t,d,e)-(m-m_{0})\big{\|}^{2}\right]\to\min_{\theta}$
where the interpolation path is given by:
@@
We parameterize the velocity field using a transformer architecture with bi-directional attention, motivated by the Diffusion Transformer Peebles and Xie [2022]. Specifically, our transformer implementation uses linear projection of input parameters into the embedding space. Time is encoded using a Timestep Embedder as proposed in [Peebles and Xie, 2022], which ensures proper time representation in the embedding space. Root Mean Square (RMS) normalization stabilizes learning dynamics. The activation function is $x=\text{ReLU}(x)^{2}$. Self-attention uses rotary position embeddings (RoPE), enabling the transformer to learn relative token positions and generalize to sequences longer than those seen during training.
The architecture varies slightly across tasks to accommodate different input data representations. Specific implementations for tasks from Section 4 are detailed in Figure 6 in Appendix B.
@@
We need to be able to predict the model parameters from different number of observation points. As mentioned before, we generate datasets with varying numbers of observation points, where each batch corresponds to samples with a specific number of points. During training, the model processes batches with different numbers of points sequentially. We propose two strategies: the first follows Algorithm 1, computing gradients for each batch and updating after a fixed number of batches.
@@ 4.2 SEIR disease model
The SEIR (Susceptible-Exposed-Infected-Removed) model is a mathematical framework used to simulate the spread of infectious diseases. In this case study, we simulate a realistic scenario where we measure the number of infected and deceased individuals at random times and use this information to recover the control parameters of the ODE system.
@@ 4.3 Permeability field inversion
Our results show that we can effectively recover the PDE coefficient using just a few strategically placed measurement points. Figure 4 demonstrates that with 8 relatively uniformly distributed points over the solution field, we can obtain an almost identical solution (approximately 2.75% relative error). The ensemble-generated $\log\kappa$ represents the mean of 50 parameter predictions from the transformer’s inference.
@@ 5 Results and Discussions
The true solution of the ODE system and the reconstructed parameter distribution, obtained using only four observation points, are illustrated in Figure 4.
@@
Figure 4: PDE coefficient and solution: true (left) and reconstructed using Flow Matching (right)
@@
We compare our method against the Metropolis-Hastings MCMC (MH-MCMC) algorithm, running it with sufficient iterations to match the error levels shown in Table 2. The results for the SEIR problem are presented in Table 3.
@@
Comparing Tables 2 and 3, we observe that MCMC requires significantly more iterations to achieve similar accuracy as Conditional Flow Matching. This computational cost is particularly significant for elliptic PDEs. Even after 10,000 MCMC iterations, the error in solving the PDE remains above 30% for 6 or more observations, while Conditional Flow Matching achieves 2-8% error in these cases. The computational time difference is substantial: MCMC takes approximately 37 minutes for 10,000 samples, while Flow Matching requires only 1.08 seconds (on CPU) per inference using Algorithm 2 for the permeability inversion problem. For the SEIR model, the CFM transformer takes approximately 0.22 seconds.
Our experiments highlight several key findings:
@@
Additionally, Conditional Flow Matching shows promise for determining optimal experiment design parameters $e$, which could further enhance its applicability to practical scientific applications.
@@ 6 Limitations and Future Work
The CFM approach is not the only generative modelling technique available. Alternative generative modeling techniques include normalizing flows, tensor methodsKoval et al. [2024], generative adversarial networks. We also need to establish the efficiency of the approach for high-dimensional Bayesian inverse problems, which will require modifications of the architecture and additional scaling of the datasets. CFM learns to sample from the distribution, whereas the Bayesian optimal experiment design requires the evaluation of the log-likelihood. Although it is in principle possible, we did not study the complexity and accuracy of the evaluation of the logarithm of the posterior distribution.
Another line of work is to introduce the guidance to the CFM objective that is enforced by the forward model. Once we have obtained the estimate of the parameter $m$, we can check if it really fits the observations; a good question is how to modify the inference procedure to correct for possible errors.
An important limitation is the study of the actual properties of the learned distribution. For the cases, when the model parameters are defined by the observations, the velocity field $v_{t}$ will likely not depend on the noise, but only on $d,e$, effectively solving the regression problem of predicting $m_{i}$ from $d_{i},e_{i}$. The usefulness and emergence of randomization with respect to the noise in CFM still needs to be studied. There are two options for using it. First, we can sample different noises, get estimates of the model parameters and plot the distributions (see the Section on the SEIR model, where some of the parameters show greater variability). The second option is to pick samples that provide better fit to the data. This still needs to be studied, since in some of the experiments we found differences in the distribution provided by MCMC and CFM. The question has to be studied in more details.
7 Code and Availability
Technical training details (architectures, learning rates, etc.) are given in Appendix B. The code is written using PyTorch framework and is publicly available at